Grade 12th passMechanics

in a straight line motion , a particle starting with velocity 2m/s covers a distance 4m in one second. Assuming uniformly accelarated motion , the final speed of the particle just after this one second is?

suman bhanja
8 Years agoGrade 12th pass

1 Answer

IHate Ppts
8 Years ago
S=ut+1/2at^2.4=2(1)+1/2a(1)^2.A=3/2.V=u+at.V=2+3/2(1).Final velocity is 3.5 m/s. Hope this helps.